# SceneSplat: Gaussian Splatting-based Scene Understanding with Vision-Language Pretraining

### v1

@@ Abstract
Recognizing arbitrary or previously unseen categories is essential for comprehensive real-world 3D scene understanding. Currently, all existing methods rely on 2D or textual modalities during training, or together at inference. This highlights a clear absence of a model capable of processing 3D data alone for learning semantics end-to-end, along with the necessary data to train such a model. Meanwhile, 3D Gaussian Splatting (3DGS) has emerged as the de facto standard for 3D scene representation across various vision tasks. However, effectively integrating semantic reasoning into 3DGS in a generalizable fashion remains an open challenge. To address these limitations we introduce SceneSplat in Fig. 1, to our knowledge the first large-scale 3D indoor scene understanding approach that operates natively on 3DGS. Furthermore, we propose a self-supervised learning scheme that unlocks rich 3D feature learning from unlabeled scenes. In order to power the proposed methods, we introduce SceneSplat-7K, the first large-scale 3DGS dataset for indoor scenes, comprising of 6868 scenes derived from 7 established datasets like ScanNet, Matterport3D, etc. Generating SceneSplat-7K required computational resources equivalent to 119 GPU-days on an L4 GPU, enabling standardized benchmarking for 3DGS-based reasoning for indoor scenes. Our exhaustive experiments on SceneSplat-7K demonstrate the significant benefit of the proposed methods over the established baselines. Our code, model, and datasets will be released at SceneSplat.
@@ 1 Introduction
The ability to interpret arbitrary queries, rather than being limited to a closed-set of categories, is crucial for 3D understanding models to generalize across diverse real-world scenarios. Traditional 3D vision systems are typically trained on fixed closed-set category labels drawn from datasets like ScanNet [5]. Such label spaces fail to capture the diversity of concepts in real environments. This gap has spurred significant interest in open-vocabulary 3D scene understanding, which aims to recognize arbitrary or unseen categories beyond a predefined taxonomy [8, 54, 15]. Achieving this capability would empower the model to reason about novel objects in a scene via natural language descriptors.
@@
The field of 3D representation is rapidly evolving. While classical 3D networks rely on point clouds or voxel grids [55, 50, 51], recent approaches inspired by radiance fields have greatly improved scene representation. Notably, 3D Gaussian Splatting (3DGS) [16] has emerged as an efficient representation, achieving state-of-the-art performance in view synthesis and geometry modeling [58, 61]. Unlike discrete point clouds, 3DGS provides a compact formulation that can be optimized per scene for 2D synthesis, while the underlying Gaussians also softly encode rich 3D structures (positions, shapes, and opacities of the regions). This makes it a unique candidate for 3D scene cues, as it naturally fuses geometry and appearance information. However, integrating semantic reasoning into 3DGS is non-trivial. The naïve way of optimizing additional semantic features in 3DGS [17] is inefficient and limited to a single scene. Consequently, generalizable open-vocabulary understanding in 3DGS remains unexplored, as no existing model directly processes 3D Gaussian parameters. Moreover, no large-scale scene-level 3DGS dataset is available to train such models, further limiting progress in this direction.
We address these limitations with SceneSplat, to our knowledge the first large-scale 3D indoor scene understanding approach that operates natively on 3D Gaussian splats. The proposed model is powered by SceneSplat-7K, a carefully curated 3DGS indoor scene dataset spanning around 7K scenes. SceneSplat introduces a 3DGS encoder that takes as input the parameters of a Gaussian-splat scene (center, scale, color, opacity) and outputs semantic features in a per-primitive manner, in a single forward pass. We leverage supervision from vision-language models to train SceneSplat’s encoder to produce CLIP-aligned embeddings, effectively bridging language and 3D without an explicit 2D fusion at runtime. Furthermore, we propose GaussSSL, a self-supervised learning scheme that unlocks rich 3D feature learning from unlabeled scenes. GaussSSL operates through three synergistic strategies, Masked Gaussian Modeling (MGM), Self-Distillation, and optionally Language-Gaussian Feature Alignment, allowing the model to separate high-level semantic signals from raw parameters alone. This self-supervised pre-training is fueled by large amounts of 3DGS scenes in the SceneSplat-7K dataset. Our contributions can be summarized as follows:
$\bullet$ We present SceneSplat-7K, a high-quality large-scale Gaussian splats dataset spanning 7K indoor scenes, which boost 3DGS scene understanding research.
$\bullet$ We propose SceneSplat, the model that unlocks open-vocabulary recognition for 3D Gaussian splats and achieves state-of-the-art zero-shot semantic segmentation performance on three fine-grained benchmarks.
$\bullet$ We incorporate annotation-free self-supervised training mechanisms on this large-scale indoor dataset and demonstrate their effectiveness in downstream indoor segmentation task.
@@ 2 Related Work
Raw Scenes 1513 380 1006 18 461 1482(scans) 1661 2194 8715 $\sim$9k
GS Scenes 1513 281 956 18 448 632(scans) 938 1982 6868 $\sim$7k
RGB Frames 2.5M 228K 1.1M 36K 77K 156K 450K 194K 4.74M
Storage 1.1TB 128GB 95GB 14GB 236GB 439GB 159GB 95GB 2.27TB
PSNR 29.07 29.49 29.11 41.25 25.93 27.46 25.65 28.64 28.17
Depth Loss 0.031 0.019 0.015 0.002 0.228 0.018 0.0136 0.122 0.064
SSIM 0.869 0.924 0.933 0.980 0.894 0.881 0.841 0.900 0.889
LPIPS 0.236 0.133 0.116 0.0396 0.157 0.335 0.380 0.208 0.230
GS per scene 1.50M 1.89M 1.89M 1.50M 2.84M 1.50M 1.52M 0.50M 1.37M
Total GS 2,269.5M 528.9M 1810.3M 23.95M 1,237.5M 948.0M 343.9M 991.0M 8.15B
GPU Time (L4) 593h 177h 594h 8h 176h 576h 290h 443h 2857h
Table 1: Dataset Statistics. The proposed SceneSplat-7K dataset includes various 3D Gaussian Splatting datasets generated from Scannet[5], Scannet++[57], Scannet++v2, Replica[45], Hypersim[42], 3RScan[48], arkitscenes[1], and matterport3d[2]. The dataset contains 6,868 scenes and 8.15 Billion 3DGS. Constructing this dataset required computational resources equivalent to 119 GPU-days on one NVIDIA L4 GPU. SceneSplat-7K achieves high-fidelity reconstruction quality with an average PSNR 28.17 dB.
@@
3D Representation Learning. Representation learning in 3D, akin to its success in the 2D image domain, is crucial for extracting meaningful features from 3D data. Previous approaches have focused on either architectural advancements or learning paradigms. Many methods have sought to map 3D data into 2D patches, leveraging conventional 2D CNNs or Vision Transformers [47, 9, 40]. Although promising, this approach often overlooks the inherent properties of 3D data due to the embedding strategy. To address this, research has shifted toward specialized architectures (i.e., PointCNN [22], PointNet [34], and Point Transformer [65, 50, 51], Pointmamba [24]). However, labeling 3D data requires extensive effort and domain expertise, making it quite challenging. This has highlighted the importance of self-supervised learning (SSL) for 3D data, with methods typically falling into contrastive or generative categories. Contrastive methods aim to differentiate similar and dissimilar examples, but often suffer from issues such as overfitting and mode collapse, exacerbated by the sparse nature of 3D data [35, 41]. On the other hand, generative methods, inspired by BERT’s mask-and-reconstruct strategy [6], have proven more effective. Masked Autoencoders [11], originally designed for 2D images, have been adapted for 3D data types, including point clouds [32, 63], meshes [25], voxels [12], and Gaussians [28], allowing the reconstruction of masked regions and facilitating the learning of robust spatial and semantic features. In this work, we propose a novel approach for language-aware 3DGS pretraining, as well as a self-supervised learning scheme that operates via Gaussian masking and self-distillation.
@@ 3 SceneSplat Dataset
We introduce SceneSplat-7K – a carefully curated dataset of 3D Gaussian Splats representing indoor scenes. The main goal was obtaining a dataset to facilitate generalizable 3DGS indoor scene understanding. The dataset contains about seven thousand scenes, including both real-world and synthetic environments. The comprehensive statistics of the introduced dataset are presented in Tab. 1. The SceneSplat-7K dataset will be publicly released, with additional details (license) available in the supplement.
@@ 3.1 Data Processing
Multiple measures are applied before, during, and after the 3DGS optimization to guarantee high-quality 3DGS scenes. Starting with the training views, we select scenes with at least 400 frames to ensure sufficient multi-view coverage. We remove blurry frames by using the variance of the Laplacian as a sharpness metric. We use gsplat [56] for 3DGS optimization. For scenes with available depth input, we apply depth loss to achieve better geometry modeling. To efficiently compress the 3DGS scene, we employ a Markov Chain Monte Carlo strategy [18] and add opacity and scale regularization. Once optimized, we filter 3DGS scenes based on the PSNR metric before using them as input for our pretraining. We refer to the supplement for per-dataset processing details.
@@ 3.2 Data Statistic
SceneSplat-7K dataset includes various 3D Gaussian Splatting datasets generated from Scannet[5], Scannet++[57], Scannet++v2, Replica[45], Hypersim[42], 3RScan[48], Arkitscenes[1], and Matterport3D[2], comprising approximately 9,000 raw scenes. Scenesplat-7K contains 6,868 processed Gaussian splatting scenes, with an average of 1.37 Million 3D Gaussians per scene and 8.15 Billion Gaussians in total, from which 4,114 high-quality Gaussian splatting scenes are selected for pertaining. Spanning a total of 4.74 Million RGB frames, SceneSplat-7K achieves high-fidelity appearance and competitive reconstruction quality, having an average PSNR of 28.17 dB, average depth loss of 0.064 m, average SSIM of 0.889, and average LPIPS of 0.230. Constructing this dataset required the equivalent of 119 days of computation on an NVIDIA L4 GPU.
@@ 4 Methodology
Building upon the SceneSplat-7K dataset, we carry out both vision-language 3DGS pretraining which enables open-vocabulary scene understanding, and self-supervised pretraining which regularizes the latent space during 3DGS parameter encoding as shown in Fig. 2. For vision-language pretraining, we first need to collect the primitive-level language labels for the 3DGS scenes (Sec. 4.1). The SceneSplat model then learns to predict vision-language features robustly from Gaussian parameters and its surroundings (Sec. 4.2). Furthermore, for self-supervised pretraining, we employ a multi-objective self-supervised training framework, integrating reconstruction and self-distillation alignment (Sec. 4.3).
@@
Unlike methods that align 3D primitives with text embeddings [8, 21] or use visual captioning [26, 59], we directly align Gaussians with the image embedding space of vision-language models (VLM), preserving richer latent semantic information. Our approach also avoids scene-specific compression methods [36, 44] that limit scalability and generalization.
As outlined in Alg. 1, we employ SAMv2 [39] for object-level segmentation and SigLIP2 [46] for feature extraction. We then use Occam’s LGS [4] to efficiently lift these 2D feature maps to a 3D Gaussian feature field in a optimization-free manner. This results in a comprehensive collection of 3DGS-feature pairs $\{(G_{i},F_{i})\}_{i=1}^{N}$ across multiple datasets, providing a solid foundation for our vision-language pretraining.
@@ 4.2 Vision-Language 3DGS Pretraining
We first adapt the transformer encoder-decoder backbone from [51] to efficiently predict high-dimensional per-primitive features corresponding to collected 3DGS language labels. More specifically, our model $g(\cdot)$, parametrized by $\theta$, maps input Gaussians to their language features:
@@
We apply three training objectives for supervision. Cosine similarity loss minimizes the angular difference between the predicted and ground truth language labels:
@@
Lastly, we use an aggregated contrastive loss to encourage separation of class-level features. Instead of contrasting every Gaussian feature individually (which would be computationally prohibitive in large scenes), we apply class-wise mean pooling. For each semantic class $c$ with sufficiently many Gaussians, we randomly split its Gaussians into two disjoint sets $\mathcal{G}_{c}^{A}$ and $\mathcal{G}_{c}^{B}$, and compute the pooled features:
@@
We then apply a bidirectional contrastive loss. First, we normalize $\bar{F}_{c}^{A}$ and $\bar{F}_{c}^{B}$ to unit length. Let $Z^{A}=\bar{F}^{A}(\bar{F}^{B})^{\top}/\tau$ and $Z^{B}=\bar{F}^{B}(\bar{F}^{A})^{\top}/\tau$, where each row in $\bar{F}^{A}$ or $\bar{F}^{B}$ corresponds to a different semantic class. The diagonal elements of $Z^{A}$ and $Z^{B}$ are the positive matches. Hence we compute a cross-entropy loss in both directions:

### v2

@@ Abstract
Recognizing arbitrary or previously unseen categories is essential for comprehensive real-world 3D scene understanding. Currently, all existing methods rely on 2D or textual modalities during training or together at inference. This highlights the clear absence of a model capable of processing 3D data alone for learning semantics end-to-end, along with the necessary data to train such a model. Meanwhile, 3D Gaussian Splatting (3DGS) has emerged as the de facto standard for 3D scene representation across various vision tasks. However, effectively integrating semantic reasoning into 3DGS in a generalizable manner remains an open challenge. To address these limitations, we introduce SceneSplat in Fig. 1, to our knowledge the first large-scale 3D indoor scene understanding approach that operates natively on 3DGS. Furthermore, we propose a self-supervised learning scheme that unlocks rich 3D feature learning from unlabeled scenes. To power the proposed methods, we introduce SceneSplat-7K, the first large-scale 3DGS dataset for indoor scenes, comprising 7916 scenes derived from seven established datasets, such as ScanNet and Matterport3D. Generating SceneSplat-7K required computational resources equivalent to 150 GPU days on an L4 GPU, enabling standardized benchmarking for 3DGS-based reasoning for indoor scenes. Our exhaustive experiments on SceneSplat-7K demonstrate the significant benefit of the proposed method over the established baselines. Our code, model, and datasets will be released at SceneSplat.
@@ 1 Introduction
The ability to interpret arbitrary queries rather than being limited to a closed set of categories is crucial for 3D understanding models to generalize across diverse real-world scenarios. Traditional 3D vision systems are typically trained on fixed closed-set category labels drawn from datasets like ScanNet [5]. Such label spaces fail to capture the diversity of concepts in real-world environments. This gap has spurred significant interest in open-vocabulary 3D scene understanding, which aims to recognize arbitrary or unseen categories beyond a predefined taxonomy [8, 54, 15]. Achieving this capability would empower the model to reason about novel objects in a scene using natural language descriptors.
@@
The field of 3D representation is rapidly evolving. While classical 3D networks rely on point clouds or voxel grids [55, 50, 51], recent approaches inspired by radiance fields have greatly improved scene representation. Notably, 3D Gaussian Splatting (3DGS) [16] has emerged as an efficient representation, achieving state-of-the-art performance in view synthesis and geometry modeling [58, 61]. Unlike discrete point clouds, 3DGS provides a compact formulation that can be optimized per scene for 2D synthesis, while the underlying Gaussians also softly encode rich 3D structures (positions, shapes, and opacities of the regions). This makes it a unique candidate for 3D scene cues, as it naturally fuses geometry and appearance information. However, integrating semantic reasoning into 3DGS is nontrivial. The naïve way of optimizing additional semantic features in 3DGS [17] is inefficient and limited to a single scene. Consequently, generalizable open-vocabulary understanding in 3DGS remains unexplored, as no existing model directly processes 3D Gaussian parameters. Moreover, no large-scale scene-level 3DGS dataset is available to train such models, further limiting the progress in this direction.
We address these limitations with SceneSplat, to our knowledge the first large-scale 3D indoor scene understanding approach that operates natively on 3D Gaussian splats. The proposed model is powered by SceneSplat-7K, a carefully curated 3DGS indoor scene dataset spanning around 7K scenes. SceneSplat introduces a 3DGS encoder that takes as input the parameters of a Gaussian-splat scene (center, scale, color, opacity) and outputs semantic features in a per-primitive manner, in a single forward pass. We leverage supervision from vision-language models to train SceneSplat’s encoder to produce CLIP-aligned embeddings, effectively bridging language and 3D without explicit 2D fusion at runtime. Furthermore, we propose GaussSSL, a self-supervised learning scheme that unlocks rich 3D feature learning from unlabeled scenes. GaussSSL operates through three synergistic strategies: Masked Gaussian Modeling (MGM), self-distillation, and optional Language-Gaussian Feature Alignment, allowing the model to separate high-level semantic signals from raw parameters alone. This self-supervised pretraining is fueled by large amounts of 3DGS scenes in the SceneSplat-7K dataset. Our contributions can be summarized as follows:
$\bullet$ We present SceneSplat-7K, a high-quality large-scale Gaussian splats dataset spanning 7K indoor scenes, which boosts 3DGS scene understanding research.
$\bullet$ We propose SceneSplat, a model that unlocks open-vocabulary recognition for 3D Gaussian splats and achieves state-of-the-art zero-shot semantic segmentation performance on three fine-grained benchmarks.
$\bullet$ We incorporate annotation-free self-supervised training mechanisms on this large-scale indoor dataset and demonstrate their effectiveness in the downstream indoor segmentation task.
@@ 2 Related Work
Raw Scenes 1613 380 1006 8 461 1482(scans) 1970 2194(regions) 9114
GS Scenes 1613 330 956 8 448 632(scans) 1947 1982(regions) 7916
RGB Frames 2.5M 228K 1.1M 16K 77K 156K 450K 194K 4.72M
Storage 600GB 152GB 447GB 7GB 251GB 235GB 577GB 492GB 2.76TB
PSNR 29.07 29.49 29.11 41.25 25.93 27.46 29.18 32.34 29.64
Depth Loss 0.031 0.019 0.015 0.002 0.228 0.018 0.0131 0.033 0.035
SSIM 0.869 0.924 0.933 0.980 0.894 0.881 0.885 0.916 0.897
LPIPS 0.236 0.133 0.116 0.0396 0.157 0.335 0.294 0.145 0.212
GS per scene 1.50M 1.56M 1.89M 1.50M 2.84M 1.50M 1.19M 1.0M 1.42M
Total GS 2419.5M 513.4M 1810.3M 12.0M 1,237.5M 948.0M 2,316.9M 1,982M 11.27B
GPU Time (L4) 593h 177h 594h 4h 176h 576h 811h 661h 3592 h
Table 1: Dataset Statistics. The proposed SceneSplat-7K dataset includes various 3D Gaussian Splatting datasets generated from ScanNet [5], ScanNet++ [57], ScanNet++ v2, Replica[45], Hypersim[42], 3RScan[48], ARKitScenes[1], and Matterport3D[2]. The dataset contains 7,916 scenes and 11.27 Billion 3DGS. Constructing this dataset required computational resources equivalent to 150 GPU-days on one NVIDIA L4 GPU. SceneSplat-7K achieves high-fidelity reconstruction quality with an average PSNR 29.64 dB.
@@
3D Representation Learning. Representation learning in 3D, akin to its success in the 2D image domain, is crucial for extracting meaningful features from 3D data. Previous approaches have focused on either architectural advancements or learning paradigms. Many methods have sought to map 3D data into 2D patches, leveraging conventional 2D CNNs or Vision Transformers [47, 9, 40]. Although promising, this approach often overlooks the inherent properties of 3D data due to the embedding strategy. To address this, research has shifted toward specialized architectures (i.e., PointCNN [22], PointNet [34], Point Transformer [65, 50, 51], and Pointmamba [24]). However, labeling 3D data requires extensive effort and domain expertise, making it quite challenging. This has highlighted the importance of self-supervised learning (SSL) for 3D data, with methods typically falling into contrastive or generative categories. Contrastive methods aim to differentiate similar and dissimilar examples, but often suffer from issues such as overfitting and mode collapse, exacerbated by the sparse nature of 3D data [35, 41]. On the other hand, generative methods, inspired by BERT’s mask-and-reconstruct strategy [6], have proven more effective. Masked Autoencoders [11], originally designed for 2D images, have been adapted for 3D data types, including point clouds [32, 63], meshes [25], voxels [12], and Gaussians [28], allowing the reconstruction of masked regions and facilitating the learning of robust spatial and semantic features. In this work, we propose a novel approach for vision-language 3DGS pretraining, as well as a self-supervised learning scheme that operates via Gaussian masking and self-distillation.
@@ 3 SceneSplat Dataset
We introduce SceneSplat-7K – a carefully curated dataset of 3D Gaussian Splats representing indoor scenes. The main goal was to obtain a dataset to facilitate generalizable 3DGS indoor scene understanding. The dataset contains about seven thousand scenes, including both real-world and synthetic environments. The comprehensive statistics of the introduced dataset are presented in Tab. 1. The SceneSplat-7K dataset will be publicly released, with additional details (licenses) available in the supplement.
@@ 3.1 Data Processing
Multiple measures are applied before, during, and after the 3DGS optimization to guarantee high-quality 3DGS scenes. Starting with the training views, we select scenes with at least 400 frames to ensure sufficient multi-view coverage. We remove blurry frames by using the variance of the Laplacian as a sharpness metric. We use gsplat [56] for 3DGS optimization. For scenes with available depth input, we apply depth loss to achieve better geometry modeling. To efficiently compress the 3DGS scene, we employ a Markov Chain Monte Carlo strategy [18] and add opacity and scale regularization. Once optimized, we filtered 3DGS scenes based on the PSNR metric before using them as inputs for our pretraining. We refer to the supplement for per-dataset processing details.
@@ 3.2 Data Statistic
SceneSplat-7K dataset includes various 3D Gaussian Splatting datasets generated from ScanNet [5], ScanNet++ [57], ScanNet++ v2, Replica [45], Hypersim [42], 3RScan [48], ARKitScenes [1], and Matterport3D [2], comprising approximately 9,000 raw scenes. SceneSplat-7K contains 7,916 processed Gaussian splatting scenes, with an average of 1.42 Million 3D Gaussians per scene and 11.27 Billion Gaussians in total, from which 4,114 high-quality Gaussian splatting scenes are selected for pretraining. Spanning a total of 4.72 Million RGB frames, SceneSplat-7K achieves high-fidelity appearance and competitive reconstruction quality, having an average PSNR of 29.64 dB, average depth loss of 0.035 m, average SSIM of 0.897, and average LPIPS of 0.212. Constructing this dataset required an equivalent of 150 days of computation on an NVIDIA L4 GPU.
@@ 4 Methodology
Building upon the SceneSplat-7K dataset, we carry out both vision-language 3DGS pretraining, which enables open-vocabulary scene understanding, and self-supervised pretraining, which regularizes the latent space during 3DGS parameter encoding, as shown in Fig. 2. For vision-language pretraining, we first need to collect primitive-level language labels for 3DGS scenes (Sec. 4.1). The SceneSplat model then learns to robustly predict vision-language features from the Gaussian parameters and their surroundings (Sec. 4.2). Furthermore, for self-supervised pretraining, we employ a multi-objective self-supervised training framework that integrates reconstruction and self-distillation alignment (Sec. 4.3).
@@
Unlike methods that align 3D primitives with text embeddings [8, 21] or use visual captioning [26, 59], we directly align Gaussians with the image embedding space of vision-language models (VLM), preserving richer latent semantic information. Our approach also avoids scene-specific compression [36, 44] which limits scalability and generalization.
As outlined in Alg. 1, we employ SAMv2 [39] for object-level segmentation and SigLIP2 [46] for feature extraction. We then use Occam’s LGS [4] to efficiently lift these 2D feature maps to a 3D Gaussian feature field in an optimization-free manner. This results in a comprehensive collection of 3DGS-feature pairs $\{(G_{i},F_{i})\}_{i=1}^{N}$ across multiple datasets, providing a solid foundation for our vision-language pretraining.
@@ 4.2 Vision-Language 3DGS Pretraining
We first adapt the transformer encoder-decoder backbone from [51] to efficiently predict high-dimensional per-primitive features corresponding to collected 3DGS language labels. More specifically, our model $g(\cdot)$, parameterized by $\theta$, maps the input Gaussians to their language features:
@@
We apply three training objectives for supervision. The cosine similarity loss minimizes the angular difference between the predicted and ground truth language labels:
@@
Lastly, we use an aggregated contrastive loss to encourage the separation of class-level features. Instead of contrasting every Gaussian feature individually (which would be computationally prohibitive in large scenes), we apply class-wise mean pooling. For each semantic class $c$ with sufficiently many Gaussians, we randomly split its Gaussians into two disjoint sets $\mathcal{G}_{c}^{A}$ and $\mathcal{G}_{c}^{B}$, and compute the pooled features:
@@
We then apply a bidirectional contrastive loss. First, we normalize $\bar{F}_{c}^{A}$ and $\bar{F}_{c}^{B}$ to unit length. Let $Z^{A}=\bar{F}^{A}(\bar{F}^{B})^{\top}/\tau$ and $Z^{B}=\bar{F}^{B}(\bar{F}^{A})^{\top}/\tau$, where each row in $\bar{F}^{A}$ or $\bar{F}^{B}$ corresponds to a different semantic class. The diagonal elements of $Z^{A}$ and $Z^{B}$ are positive matches. Hence, we compute a cross-entropy loss in both directions: